OpenStudy (anonymous):

How do I solve this: ln(x+1)-ln(x-1)=3

OpenStudy (anonymous):

\[\begin{align} ln\left(\frac{x+1}{x-1}\right)&=3\\ \frac{x+1}{x-1}&=e^3\\ x+1&=e^3 x - e^3\\ e^3+1&=(e^3-1)x\\ x&=\frac{(e^3+1)}{(e^3-1)}\\ \end{align}\]

OpenStudy (anonymous):

Thank you!